OpenStudy (anonymous):

Divide the following using synthetic division: x2-2 x2+2 x3+2 x3-2

OpenStudy (anonymous):

\[x ^{3}+3x ^{2}-2x-6 this is the equation

OpenStudy (amistre64):

synthetic strips off the variables and fills in the missing parts with 0s is the key to remember

OpenStudy (amistre64):

you have provided one equation; what is it being divided by?

OpenStudy (amistre64):

1x^3 +3x^2 -2x -6 1 3 -2 -6 0 ------------- d| 1 this is the best i got for starters, we are missing a part for the d tho

OpenStudy (anonymous):

sorry the equation is over x + 3

OpenStudy (amistre64):

my format is off at the moment; someone touched my stuff :)

OpenStudy (amistre64):

x+3 = 0 when x=-3; so we use d=-3 if that had formated right

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

q w e

OpenStudy (amistre64):

formats still off

OpenStudy (anonymous):

thats ok

OpenStudy (akshay_budhkar):

how was ur dance class?

OpenStudy (amistre64):

it was ok, we talked about the teach and cried alot :)

OpenStudy (akshay_budhkar):

cried ???

OpenStudy (akshay_budhkar):

Was it a reunion?

OpenStudy (amistre64):

1 3 -2 -6 / -3 = 1+0=1,*-3=-3, +3=0,*-3 = 0, -2 = -2, * -3 = 6, - 6 = 0

OpenStudy (amistre64):

its a factor if you can follow the line :) and it produces: x^2 -2 as a result

OpenStudy (anonymous):

alright thanks for you help :)

OpenStudy (amistre64):

x + 3 x^2 -2 ------ x^3 +3x^2 -2x -6 ------------- x^3 +3x^2 -2x -6 was the original right? and thats spose to be on seperate lines tho

OpenStudy (anonymous):

yup thanks

OpenStudy (amistre64):

good luck with it all :) i gotta run to dance class #2

OpenStudy (akshay_budhkar):

lol